### v1

# HIRO: Hierarchical Information Retrieval Optimization

Abstract
Large Language Models (LLMs) excel in natural language tasks but face limitations due to static training datasets, resulting in outdated or contextually shallow responses. Retrieval-Augmented Generation (RAG) addresses this by integrating real-time external knowledge, enhancing model accuracy and credibility, especially for knowledge-intensive tasks. However, RAG-enhanced LLMs struggle with long contexts, causing them to ”choke” on information overload, compromising response quality. Recent RAG applications use hierarchical data structures for storing documents, organized at various levels of summarization and information density. In this context, we introduce HIRO (Hierarchical Information Retrieval Optimization), a novel querying approach for RAG applications using hierarchical structures for storing documents. HIRO employs DFS-based recursive similarity score calculation and branch pruning to minimize the context returned to the LLM without informational loss. HIRO outperforms existing querying mechanisms on the NarrativeQA dataset by an absolute performance gain of 10.85%.
Figure 1: Architectural Overview of a Retrieval-Augmented Language Model (RALM). A query is inputted, processed through an embedding model to generate a vector representation, which is then matched in a vector database to retrieve relevant contexts. These contexts are fed, along with the original query, into the LLM, resulting in an informed response.
Introduction
The advent of Large Language Models (LLMs) has brought about a significant transformation in artificial intelligence and natural language processing, enabling significant advancements in various tasks such as text generation, translation, and question-answering (Bubeck et al. 2023). With growing size and complexity, these models have evolved into highly effective standalone knowledge repositories, embedding vast amounts of factual information within their parameters (Petroni et al. 2019; Talmor et al. 2020). However, LLMs are inherently constrained by the static nature of their training datasets (Minaee, Mani, and Aggarwal 2024), lacking domain-specific knowledge and the adaptability to continuously evolving real-world information where the breadth and depth of knowledge required far exceed the capacity of any static training corpus. Fine-tuning, though used to tailor models, falls short in effectively integrating evolving data, especially when confronted with the need to continuously update and incorporate domain-specific or real-time information (Lewis et al. 2020). This challenge underscores a critical limitation in LLM design.
In response to these limitations, the field has witnessed the emergence of Retrieval-Augmented Generation (RAG) models (Lewis et al. 2020), leading to the development of Retrieval-Augmented Language Models (RALMs). These paradigms enhance LLMs by integrating dynamic external knowledge through retrieval mechanisms. Inspired by the success of open-domain question answering systems, these approaches utilize indexed text databases, enhancing model responses by presenting retrieved information alongside input queries. RALMs not only boost the models’ ability to provide current, domain-specific knowledge but also introduce interpretability and source traceability, contrasting with the opaque parametric knowledge of traditional LLMs (Ram et al. 2023; Akyürek et al. 2022). This represents a significant stride towards overcoming the inherent limitations of static knowledge systems.
Amidst the evolving landscape of information retrieval for natural language processing, novel approaches to information retrieval, such as the use of hierarchical data structures for organizing documents (Kanagavalli and Maheeja 2016; Wang et al. 2024; He et al. 2024; Sarthi et al. 2024), represents a significant step forward. Techniques like RAPTOR: Recursive Abstractive Processing for Tree-Organized Retrieval (Sarthi et al. 2024), which clusters and summarizes texts into a tree structure, showcases the potential for such methodologies to overcome traditional limitations by offering both granular and high-level insights. Yet, the adoption of hierarchical data structures presents challenges, notably the overwhelming amount of context they can return.
These hierarchical data structures typically employ fixed strategies like returning a predefined top-$k$ most relevant nodes following Tree Traversal and Collapsed Tree Querying (Sarthi et al. 2024). This approach results in a static quantity of data being retrieved, which may not align with the requirements of the query. For queries demanding extensive information, this can result in too little data being returned, whereas for simpler inquiries, it might overwhelm Large Language Models (LLMs) with excessive context. This static definition of ’k’ does not account for the variability in information needs, potentially leading to an information choke in LLMs. Overloading LLMs with too much context can degrade their ability to generate coherent and relevant responses, particularly in handling large contexts (Liu et al. 2024; Sun et al. 2021; Levy, Jacoby, and Goldberg 2024). This limitation hampers the models’ performance and undermines the efficiency and applicability of Retrieval-Augmented Language Models (RALMs) in practical scenarios.
To address this challenge, our research proposes a novel querying technique designed specifically for optimizing the use of hierarchical data structures within retrieval-augmented generation systems. By applying depth-first search (DFS)-based recursive similarity thresholding and branch pruning, our approach aims to selectively minimize the context delivered to LLMs. This strategy intends to mitigate the risk of information overload, ensuring that LLMs receive a distilled yet comprehensive information set, thereby enabling more accurate and efficient response generation.
Related Work
Retirieval Augmented Generation (RAG)
Retrieval-Augmented Generation (RAG) (Lewis et al. 2020), enhances Large Language Models (LLMs) by integrating real-world data from external databases, thus birthing Retrieval Augmented Language Models (RALMs). This approach ensures generated content is both contextually relevant and factually accurate, addressing the issue of LLMs producing plausible but incorrect ”hallucinated” information (Shuster et al. 2021). By enhancing the accuracy and reliability of AI-generated text, RAG heralds a new era of trustworthiness and clarity in machine-generated communications.
RAG’s widespread use in domains like information retrieval and question-answering underscores its utility. Yet, its scalability and efficiency issues, especially in handling large datasets and complex structures, highlight the crucial need for ongoing optimization research to extend its applications and enhance performance.
Adoption of Hierarchical Data Structures for Documents
The adoption of hierarchical structures in RAG models, exemplified by RAPTOR (Sarthi et al. 2024), represents a pivotal shift towards more organized and efficient information storage and retrieval. RAPTOR’s approach, through recursive summarization to create a hierarchical tree, facilitates nuanced access to information at various levels of abstraction. This innovation not only improves coherence and information density tailored to specific tasks but also encourages exploration into other hierarchical frameworks like graphs for enhanced document interrelation preservation. Such advancements signal a new phase in computational linguistics, focusing on sophisticated, hierarchical data use for better knowledge representation and retrieval.
Figure 2: Tree Traversal Querying Mechanism. The top-$k$ root nodes are selected based on their similarity to the query. In the next layer, the children of these nodes are considered, and the top-$k$ nodes are again selected based on similarity. This process continues until reaching the leaf nodes.
Figure 3: Collapsed Tree Querying Mechanism. The tree is flattened into a single layer, and the top-$k$ nodes are directly selected based on their similarity to the query embedding. The returned context consists of the most relevant nodes from this single-layer selection.
Figure 4: HIRO Querying Mechanism. Our method’s recursive process, which employs Selection and Delta Thresholds to filter and prune document graphs for optimized information retrieval. It highlights how relevant contexts are dynamically identified and retrieved based on a balance of specificity and information breadth, tailored to each unique query.
Existing Querying Mechanisms
Existing Querying Mechanisms in Retrieval-Augmented Generation (RAG) models have undergone significant transformations to better address the dynamic and complex nature of queries. These models initially employed static retrieval methods, fetching a set number of documents without considering the query’s complexity. However, advancements like the Tree Traversal and Collapsed Tree Approach have introduced more adaptive methodologies (Sarthi et al. 2024).
The Tree Traversal approach progressively selects the top-$k$ most relevant nodes from each layer of the hierarchical data based on similarity. This method allows for adjusting the depth and number of nodes to control the information’s specificity and breadth.
Conversely, the Collapsed Tree approach simplifies retrieval by flattening the hierarchy, enabling simultaneous comparison of all nodes.
Despite these innovations, a limitation arises from the static quantity of data retrieved, which may not always align with the query’s needs, leading to potential information overload for LLMs. This static $k$ value, irrespective of query complexity, can hamper LLMs’ performance by either providing insufficient data for complex queries or overwhelming them with excessive context for simpler ones.
Moreover, these methods tend to return both parent (summary of child nodes) and child nodes in hierarchical document structures due to matching similarity levels, this may not be efficient. This inefficiency arises from the duplicated information caused by overlapping content between the parent and child nodes. Consequently, optimizing the retrieval process to prevent redundant data retrieval becomes vital for improving the efficiency and effectiveness of LLMs in handling queries within hierarchical document structures.
Detailed pseudocode for the querying mechanisms discussed, including the Tree Traversal and Collapsed Tree approaches, is available in Appendix A1 for further reference.
Recursive Similarity Check and Branch Pruning for Querying
Our approach introduces Recursive Similarity Thresholding and Branch Pruning for dynamic query response optimization. Leveraging the Similarity Threshold ($S$) and Delta Threshold ($\Delta$) hyperparameters, it tailors retrieval processes to query intricacies. $S$ filters document graphs based on similarity scores, while $\Delta$ refines relevance by pruning branches. This adaptable approach enables fine-tuning of hyperparameters at document or corpus levels, catering to specific informational needs and contexts.
Model Average Context Length
RAPTOR with Collapsed Tree Querying 2253.34586466
RAPTOR with Tree Traversal Querying 2762.46115288
RAPTOR with HIRO Querying 1864.29344498
Table 1: Average Context Length (in tokens) Returned With and Without HIRO on the NarrativeQA Dataset.
Model ROUGE-L F1 BLEU-1 BLEU-4 METEOR
RAPTOR with Collapsed Tree Querying 0.096542025 0.165350602 0.084131409 0.125686168
RAPTOR with HIRO Querying 0.120835776 0.173288844 0.089863157 0.134032758
Table 2: NarrativeQA Dataset Performance Absolute Comparison. Absolute performance comparison of RAPTOR with conventional querying and HIRO querying on the NarrativeQA dataset using SBERT, with GPT-3.5 Turbo as the language model. HIRO outperforms the existing querying mechanisms across all metrics.
Model ROUGE-L F1 BLEU-1 BLEU-4 METEOR
RAPTOR with Collapsed Tree Querying 0.012579121 0.021544661 0.010962057 0.01637651
RAPTOR with HIRO Querying 0.016045879 0.023011163 0.011933 0.017798317
Table 3: NarrativeQA Dataset Efficiency-Adjusted Performance Scores. Performance scores of RAPTOR with conventional querying and HIRO querying on the NarrativeQA dataset, using SBERT and GPT-3.5 Turbo as the language model. Scores are adjusted by the logarithm of context length (in tokens) to account for efficiency. HIRO querying demonstrates superior performance across all efficiency-adjusted metrics compared to conventional querying
Model Accuracy Precision Recall F1 Score
RAPTOR with Collapsed Tree Querying 0.186201 0.476749 0.488198 0.480011
RAPTOR with HIRO Querying 0.182493 0.449404 0.444282 0.44546
Table 4: Absolute Performance Metrics on the QuALITY Dataset. Performance comparison of RAPTOR with and without HIRO Querying on the QuALITY Dataset.
Figure 5: Absolute Impact of HIRO on the NarrativeQA Dataset. Comparison of performance metrics between top-$k$ Collapsed Tree Querying and HIRO Querying with RAPTOR. The graph presents absolute scores for the ROUGE, BLEU-1, BLEU-4, and METEOR metrics.
Figure 6: Efficiency Analysis of HIRO on the NarrativeQA Dataset. Efficiency-adjusted performance of RAPTOR with HIRO Querying compared to Collapsed Tree Querying. HIRO Querying not only improves absolute performance but also demonstrates significant efficiency gains
Figure 7: (a) Confusion Matrix for RAPTOR with Collapsed Tree Querying on the QuALITY dataset; (b) Confusion Matrix for RAPTOR with HIRO Querying on the QuALITY dataset.
Methods
Our approach introduces Recursive Similarity Score Calculation and Branch Pruning to optimize query responses through dynamic retrieval processes tailored to the specifics of each query. This method is grounded in the utilization of two hyperparameters: the Selection Threshold ($S$) and the Delta Threshold ($\Delta$), which together orchestrate a dynamic retrieval process tailored to the query’s intricacies.
Selection Threshold ($S$) serves as an initial filter, identifying document graphs for further exploration based on similarity scores between the query and parent nodes. Only graphs with parent nodes surpassing this threshold proceed, ensuring focus on potentially relevant content.
Algorithm 1 HIRO Querying
Input: Query $query$, Tree $tree$, Selection Threshold $S$, Delta Threshold $\Delta$
Output: Context nodes
1:  $context\leftarrow[]$
2:  $earmarked\leftarrow[]$
[…]
4:  for node in $l_{\text{current}}$ do
5:     $score\leftarrow\text{similarity}(query,node)$
6:     if $score>S$ then
7:        $earmarked.\text{append}(node)$
8:     end if
9:  end for
10:  $context\leftarrow\text{evaluateChildren}(query,earmarked,S,\Delta)$
11:  return $context$
Algorithm 2 Evaluate Children
Input: Query $q$, Nodes $nodes$, Selection Threshold $S$, Delta Threshold $\Delta$
Output: Local context nodes
1:  $local\_context\leftarrow[]$
2:  for parent_node in $nodes$ do
3:     $parent\_similarity\leftarrow\text{similarity}(q,parent\_node)$
4:     for node in $parent\_node.\text{children}$ do
5:        $score\leftarrow\text{similarity}(q,node)$
6:        $delta\leftarrow score-parent\_similarity$
[…]
8:           $local\_context.\text{append}(node.\text{content})$
9:        else
10:           $local\_context.\text{extend}(\text{evaluateChildren}(q,[node],$ $\quad\quad S,\Delta))$
11:        end if
12:     end for
13:  end for
14:  return $local\_context$
Delta Threshold ($\Delta$) refines this by evaluating each child node within accepted graphs. Branches with similarity below $S$ are pruned, eliminating less relevant data. For remaining child nodes, a comparison of their similarity score with their parent’s determines further action. If the difference exceeds $\Delta$, indicating heightened relevance, the process recursively continues with the child node; otherwise, the parent node is retained for context.
Below, we outline the step wise functioning of this method:
  1. 1.
    Start at the root layer of the hierarchical data structure, and compute the similarity between query embedding and the embeddings of all nodes present at this initial layer
  2. 2.
    Earmark the nodes whose similarity scores surpass the predefined Selection Threshold ($S$) for further exploration
  3. 3.
    For each earmarked node, delve into its child nodes, calculating the similarity between the query embedding and each of these subsequent nodes
  4. 4.
    Compare the difference between the similarity of child node and parent node with the Delta Threshold ($\Delta$) to evaluate whether the similarity score of a child node sufficiently exceeds that of its parent node; if so, this child node is marked for further recursive evaluation. Otherwise the parent node is added to context and the remaining branch is pruned
  5. 5.
    Terminate the search if the current node is a leaf node and sufficiently exceeds the delta threshold from its parent, or when no further nodes meet the threshold criteria
  6. 6.
    The selected nodes are then aggregated, forming an optimized and coherent context for the LLM
The Selection Threshold ($S$) governs the scope of data examined, controlling the breadth of information explored. In contrast, the Delta Threshold ($\Delta$) governs the extent of information examined, managing the depth of the search.
While traversing deeper into the tree, similarity scores naturally escalate. Nonetheless, returning excessively granular context becomes impractical due to the resultant escalation in context length and computational demands. The Delta Threshold serves as the pivotal point that delineates the balance between retaining pertinent information and mitigating the risk of inundating the system with excessive detail.
Detailed pseudocode for the HIRO Querying Mechanism discussed is available in Algorithm 1 and Algorithm 2.
Experiments
Datasets Used: We measured the performance of HIRO paired with RAPTOR across two question-answering datasets - NarrativeQA and QuALITY.
NarrativeQA is an English-language dataset, that consists of stories, which are books and movie scripts, with human written questions and answers based solely on human-generated abstractive summaries. It tests comprehension with 1567 stories and 46,765 questions (Kočiský et al. 2018), focusing on narratives of 50,000-100,000 words. This dataset tests systems’ ability to prune irrelevant details for key story elements, showcasing querying optimization for narrative understanding and information retrieval. We measure performance on this dataset using the standard ROUGE, BLEU-1, BLEU-4, METEOR.
The QuALITY dataset (Question Answering with Long Input Text, Yes!) is designed for multiple-choice question answering (QA) tasks involving lengthy English texts, ranging from 2000 to 8000 tokens per article. It comprises a total of 6,737 questions, with a challenging subset, QuALITY-HARD, containing 3,360 particularly difficult questions (Pang et al. 2021). This dataset aims to evaluate the ability of QA models to process and extract relevant information from extensive text passages. Performance is measured using the standard F1 score and Accuracy, highlighting the effectiveness of various pruning methods in identifying crucial information within long documents.
Controlled Baseline Comparisons
To establish baseline performance, we utilize RAPTOR with SBERT (Reimers and Gurevych 2019) as the embedding model and GPT 3.5 Turbo (Brown et al. 2020) as the reader/language model. This configuration represents the state-of-the-art approach in retrieval-augmented generation with RAPTOR. We compare the performance of RAPTOR coupled with SBERT and GPT 3.5 Turbo with and without HIRO querying implemented. RAPTOR conventionally selects the top-$k$ nodes after flattening the tree (Collapsed Tree Querying), and we analyze the difference in performance between the conventional approach and HIRO.
Additionally, we provide insights into the token length of the final context returned to the LLM after querying, offering a deeper understanding of the tradeoff between computational resources and performance gains.
Furthermore, the performance of LLMs degrades non-linearly with prompt token length (Levy, Jacoby, and Goldberg 2024), we also present normalized performance scores for conventional querying and HIRO querying, adjusted by the logarithm of context length (in tokens) under efficiency analysis.
Experimental Setup
For each dataset, we fine-tune the values of the two hyperparameters - Selection Threshold ($S$) and Delta Threshold ($\Delta$), to optimize HIRO’s performance. Bayesian Optimization is utilized for fine-tuning, with insights provided in Appendix C.
Experimental Results
Model Performance on Sequence Generation Tasks: Table 1 shows the average context length in tokens returned by different querying mechanisms with the RAPTOR model on the NarrativeQA dataset. The HIRO querying mechanism returned shorter contexts on average compared to tree traversal and collapsed tree querying.
Model Performance on NarrativeQA Dataset: Table 2 presents absolute performance comparison of RAPTOR with collapsed tree querying vs HIRO querying on the NarrativeQA dataset using SBERT and GPT-3.5 Turbo. HIRO querying outperformed conventional querying across all metrics. Table 3 shows the efficiency-adjusted performance scores, calculated by adjusting the absolute scores by the logarithm of context length. HIRO querying demonstrated superior efficiency-adjusted performance over conventional querying on all metrics.
Figure 5 visualizes the absolute impact of HIRO querying, showing the performance gain over collapsed tree querying for ROUGE, BLEU-1, BLEU-4, and METEOR metrics. Figure 6 depicts the efficiency analysis, illustrating that HIRO querying improves both absolute performance as well as efficiency compared to collapsed tree querying.
Model Performance on the QuALITY Dataset: Table 4 compares the absolute accuracy, precision, recall and F1 score of RAPTOR with HIRO querying vs collapsed tree querying on the QuALITY dataset. HIRO querying showed slightly lower accuracy and F1 score, compared to collapsed tree querying.
Figure 7 presents the confusion matrices for RAPTOR with HIRO querying against collapsed tree querying on QuALITY dataset.
Conclusion
This research introduces Hierarchical Information Retrieval Optimization (HIRO), a novel technique that enhances Retrieval-Augmented Generation (RAG) systems using hierarchical document structures. By integrating Depth-First Search (DFS)-based recursive similarity thresholding and branch pruning, HIRO effectively manages the context provided to Large Language Models (LLMs), addressing the prevalent challenges of information overload. The approach not only improves the quality and relevance of responses but also demonstrates increased efficiency in data retrieval, reducing computational load and enhancing overall model performance.
Reproducibility Statement
Language Models used for QA and Summarization
OpenAI GPT 3.5 Turbo (Levy, Jacoby, and Goldberg 2024)https://platform.openai.com/docs/models/gpt-3-5-turbo
Datasets Used
NarrativeQA (Kočiský et al. 2018)https://github.com/nyu-mll/quality
QuALITY (Pang et al. 2021)https://github.com/google-deepmind/narrativeqa
Source Code
RAPTOR + HIRO Querying Mechanismhttps://github.com/krishgoel/hiro
Original RAPTOR Source Codehttps://github.com/parthsarthi03/raptor
Appendix A Appendix
Pseudocode for Existing Tree Querying Mechanisms
The following section outlines two distinct conventional querying methods used for querying trees - Tree Traversal Querying in Algorithm 3 and Collapsed Tree Querying in Algorithm 4.
Algorithm 3 Tree Traversal Querying
Input: Tree $tree$, Query $query$, Integer $k$
Output: Top $k$ nodes
[…]
2:  for layer in range($tree$.num_layers) do
3:     $top\_k\leftarrow[]$
4:     for node in $S_{\text{current}}$ do
5:        $score\leftarrow\text{similarity}(query,node)$
6:        $top\_k.\text{append}(node,score)$
7:     end for
8:     $S_{\text{layer}}\leftarrow\text{sorted}(top\_k)[:k].\text{nodes}$
9:     $S_{\text{current}}\leftarrow S_{\text{layer}}$
10:  end for
11:  return $\bigcup_{i=0}^{k}S_{i}$
Algorithm 4 Collapsed Tree Querying
Input: Tree $tree$, Query $query$, Integer $k$, Integer $max\_tokens$
Output: Top nodes within token limit
1:  $tree\leftarrow\text{flatten}(tree)$
2:  $top\_nodes\leftarrow[]$
3:  for node in $tree$ do
4:     $top\_nodes.\text{append}((node,\text{similarity}(query,node)))$
5:  end for
6:  $top\_nodes\leftarrow\text{sorted}(top\_nodes)$
7:  $result\leftarrow[]$
8:  $total\_tokens\leftarrow 0$
9:  for node in $top\_nodes$ do
10:     if $total\_tokens+node\text{.token\_size}<max\_tokens$ then
11:        $result.\text{append}(node)$
12:     end if
[…]
14:  end for
15:  return $result$
Bayesian Optimization for Fine-Tuning of Hyperparameters
In our efforts to optimize the performance of our model on various datasets, we focus on fine-tuning key hyperparameters. The target function for optimization is defined as the performance of the model, which we quantify through a weighted or normalized sum of various metrics relevant to each specific dataset. These metrics may include BLEU Scores, ROUGE Scores, METEOR Scores, Precision, Recall, F1 score, or other dataset-specific performance indicators. The performance function being optimized for each dataset is provided below.
We utilize Bayesian Optimization to fine-tune the hyperparameters Selection Threshold ($S$) and Delta Threshold ($\Delta$) for various datasets. Bayesian Optimization is chosen for its proficiency in optimizing complex, multi-metric objective functions (Snoek, Larochelle, and Adams 2012). Bayesian Optimization utilizes a probabilistic model to estimate the objective function landscape and an acquisition function to efficiently explore this landscape by focusing on areas likely to offer improvements.
NarrativeQA Dataset Performance Function:The performance function for the NarrativeQA dataset is designed to evaluate the model’s ability to comprehend and summarize narratives. This function incorporates the following key metrics, ensuring thorough evaluation across different dimensions of summarization quality. To maintain consistency and comparability, each metric score is normalized individually.
  • ROUGE (Recall-Oriented Understudy for Gisting Evaluation): Measures the overlap of n-grams between the generated summaries and reference summaries to assess the recall of content.
  • BLEU-1 and BLEU-4 (Bilingual Evaluation Understudy): BLEU-1 measures the match of single words (1-gram) between the generated and reference texts, focusing on lexical accuracy. BLEU-4 extends this evaluation to 4-grams, assessing the coherence and order of longer phrases in the generated text.
  • METEOR (Metric for Evaluation of Translation with Explicit ORdering): Evaluates translation QuALITY by aligning the generated text with reference texts, accounting for synonymy and stemming, and applying varied weights to different matching components.
$P=ROUGE+{BLEU}_{1}+{BLEU}_{4}+METEOR$
Normalization process ensures that each metric contributes equally to the performance function, allowing for a balanced and fair evaluation of the model’s capabilities.
QuALITY Dataset Performance Function: For the QuALITY dataset, the performance function focuses exclusively on the F1 score to assess the model’s capability in understanding and accurately answering multiple-choice reading comprehension questions. The F1 score is a balanced metric that considers both precision and recall, making it an ideal choice for evaluating the model’s overall effectiveness in this context.
$P=F1$
$\implies P=2\cdot\frac{Precision\cdot Recall}{Precision+Recall}$